velmalovesshaggy145:

A function can be factored as f(x)=(x-11)^3(x+1)^4. How many zeroes does the graph have, what are they, and does the graph touch or cross at those zeroes? -Khi

hannahj:

The multiplicity of a root is the number of times the root appears. x = 11 (Multiplicity of 3) x =−1 (Multiplicity of 4)

hannahj:

hold on Im not done

YourlocalRandom2000:

i think The graph of the function f(x) = (x-11)^3(x+1)^4 has two zeroes: x = 11 with a multiplicity of 3, and x = -1 with a multiplicity of 4. The graph crosses the x-axis at x = 11 and touches the x-axis at x = -1.

hannahj:

omg the graph wont load

hannahj:

https://www.mathway.com/Algebra if you need the graph go here and type in f(x)=(x-11)^3(x+1)^4

YourlocalRandom2000:

Or use desmos

hannahj:

@yourlocalrandom2000 wrote:
Or use desmos
yep that works too :)